OpenStudy (anonymous):

Solve this equation: –9h – 6 + 12h + 40 = 22.

OpenStudy (anonymous):

4!!! Thank you! I still don't understand these equations no matter how hard I try to understand how to do them. I am studying these equations and hoping that I will pass my test. Thank you for your help!!!!

OpenStudy (anonymous):

First, add like terms and get 3h + 34 = 22, then bring the 34 over to the other side by subtracting 34 from both sides. 3h + 34 - 34 = 22 - 34 to get 3h = -12, then divide 3 from both sides and you get 3h/3 = -12/3, h = -12/3 = ?

OpenStudy (anonymous):

Sorry, OS is lagging my messages. I just deleted one. Anyways, the answer would be -4, because the 12 is negative.

OpenStudy (anonymous):

Ok, thank you!!

OpenStudy (anonymous):

:] no problem. I hope you do well on your test.

OpenStudy (anonymous):

Thank you and so do I!!!! :-)